solve for x
3x^2 -5x -4 = x(2x +3) +5

Answer by nerdybill(7384)

3x^2 -5x -4 = x(2x +3) +5
3x^2 -5x -4 = 2x^2 +3x +5
x^2 -5x -4 = 3x +5
x^2 -8x -4 = 5
x^2 -8x -9 = 0
(x+1)(x-9) = 0
x = {-1, 9}